A student randomly and independently samples and assigns 15 participants to one of 4 conditions. She gives each group a treatment she believes will affect performance on the dependent measure. Higher scores indicate better performance. The ANOVA summary table is from the data that she has collected. Find the dfwithin. If needed, please round to two decimal places.

Source SS df MS F
Between 82
Within 31
Answer #1

Ans:

Total number of participants,N=15*4=60

(15 participants for each group)

df(treatment)=4-1=3

df(within)=60-4=56

Source SS df MS F
Between 82 3 27.33 49.38
Within 31 56 0.55
Total 113 59